A 2.50 L sample of butane gas (C4H10), measured at 22.0 oC and 1.20 atm pressure, is combusted completely and the carbon dioxide gas collected at the same pressure and temperature. What volume of CO2 is produced?
1
Expert's answer
2020-12-16T05:54:44-0500

2C4H10 + 13O2 → 8CO2 + 10H2O

"PV= nRT"

n= (1.2×2.50)/(8.314× 295.15)

n= 0.0012moles of butane

The moles of CO2 produced are (0.0006× 8)= 0.0048moles of CO2

There the volume become;

V= (0.0048 × 8.314× 295.15)/ 1.2

V=9.82L